# In celebration of Prof. Xun Yu Zhou’s 60th birthday

### v1

@@ 1 Introduction
The recent generative modeling literature has seen a surge in the use of Schrödinger bridges. In these applications, one typically starts or chooses some distribution $\mu_{\rm ini}$} that is easy to sample from, and tries to "learn" an unknown distribution $\mu_{\rm tar}$} of a given data set. By numerically approximating the solution to the Schrödinger bridge problem, one can generate unlimited samples (i.e., synthetic data points that resemble the original data set). One such algorithm is presented by De Bortoli et al. [20] and Vargas et al. [64], who approximate the iterative proportional fitting procedure (Deming-Stephan [21]), using score matching with neural networks and maximum likelihood, respectively. Concurrently, Wang et al. [66] proposed a two-stage method with an auxiliary bridge handling possible non-smooth $\mu_{\rm tar}$}. Some more recent developments include [13, 56, 49, 52, 67, 32, 63, 54, 62, 2] as well as developing optimal transport techniques for generative AI tasks [44, 5, 42, 68, 1].
@@
In this paper, in light of Garg et al. [29], we study a soft-constrained Schrödinger bridge problem (SCSBP). Mathematically, we consider a (smooth) penalty function $G:\mathscr{P}_{2}(\mathbb{R}^{d})\mapsto\mathbb{R}_{+}$, satisfying $G(\mu;{\mu_{\rm tar}})=0$, where $\mathscr{P}_{2}(\mathbb{R}^{d})$ denotes the 2-Wasserstein space on $\mathbb{R}^{d}$, and ${\mu_{\rm tar}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ is some given “target” distribution. For each $k\in\mathbb{N}$ and a given initial distribution ${\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, we consider the following stochastic control problem with dynamics:
@@
where $\mathbb{P}_{X^{\alpha}_{T}}$ is the law of $X^{\alpha}_{T}$ and the control $\alpha$ is chosen from a square-integrable progressively measurable admissible control set $\cal A$}. The goal is to find $V^{k}:=\inf_{\alpha\in{\cal A}}J^{k}(\alpha)$ and optimal control $\alpha^{k}$, for each $k\in\mathbb{N}$, and study the limiting behavior of $\{\alpha^{k}\}$ and $\{V^{k}\}$. Clearly, the dependence of the terminal cost on $\mathbb{P}_{X^{\alpha}_{T}}$ renders this relaxed formulation a non-standard stochastic control problem, leading to a McKean-Vlasov type of stochastic control. In contrast to Garg et al. [29], which focuses on the case where the penalty $G$ is given by the KL divergence and $\mu_{\rm ini}$ is a delta measure, we investigate the problem under more general cost functions and initial distributions.
Compared to existing methods using SBP under hard constraints, there are several advantages to considering SCSBP. First, when the KL divergence between $\mu_{\rm tar}$} and $\mathbb{P}_{X^{\alpha}_{T}}$ is infinite, the Schrödinger bridge does not admit a solution, whereas SCSBP always does Garg et al. [29]. More importantly, the penalty parameter $k$ acts as a regularization factor, preventing the algorithm from overfitting to $\mu_{\rm tar}$, which is crucial for certain generative modeling tasks with limited data [30, 45]. In addition, SCSBP provides a more general framework in generative AI, with applications beyond pre-training in data generation, and can be applied to fine-tuning and transfer learning (see Examples 2.6 and 2.7).
@@ 1.1 Outline of the Main Results and Contributions
The similar results in the case of a general initial distribution is much more involved. Among other things, we establish and/or extend some recently observed stability results of the solutions to the SBP, as the foundation for a fixed-point argument. A key element in our argument is the continuous dependence (or stability) of a mapping that is well known in the (static) optimal transport literature. Specifically, for any $({\mu_{\rm ini}},{\mu_{\rm tar}})\in\mathscr{P}_{2}(\mathbb{R}^{d})\times\mathscr{P}_{2}(\mathbb{R}^{d})$, it is known (cf. e.g., [7]) that there exists a unique pair of $\sigma$-finite measures $(\nu_{0},\nu):={\cal T}({\mu_{\rm ini}},{\mu_{\rm tar}})$ such that the measure $\pi\in\mathscr{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ defined by
$\pi(E)=\int_{E}p(T,y;0,x)\nu_{0}(dx)\nu(dy),\qquad E\in\mathscr{B}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ (1.1)
has the marginals $\mu_{\rm ini}$} and $\mu$, where $p(\cdot,\cdot;\cdot,\cdot)$ is the transition density of a given diffusion process. If we fix $\mu_{\rm ini}$}, and denote $\rho^{\mu}$ to be the density of the measure $\nu={\cal T}(\mu)$, $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, then it turns out that the solution to the SCSBP is the fixed-point of the mapping $\Gamma:\mathscr{P}_{2}(\mathbb{R}^{d})\mapsto\mathscr{P}_{2}(\mathbb{R}^{d})$, where
$\Gamma(\mu):={\operatorname*{arg\,min}}_{\nu\in\mathscr{P}_{2}(\mathbb{R}^{d})}\left\{kG(\nu)+\mathbb{E}_{X\sim\nu}[\log\rho^{\mu}(X)]\right\}.$ (1.2)
The successful application of Schauder’s fixed-point theorem on the Wasserstein space relies on several key elements, in particular a crucial continuous dependence result of the mapping $\mu\mapsto\rho^{\mu}$, for which we introduce an auxiliary entropic optimal transport problem, and identify its solution to the measure $\pi$ in (1.1). By utilizing some important stability results of the corresponding Schrödinger potentials (Proposition 5.8), together with some arguments in the spirit of the converse of Scheffé’s theorem (i.e., the weak convergence in Prohorov metric vs. the convergence in densities) as well as the so-called $\Gamma$-convergence of the minimizers of the optimization problem (1.2), we are able to verify the required properties so the mapping $\Gamma$ has a fixed-point (Theorem 6.1). As a direct consequence, we establish that the optimal control of the SCSBP converges linearly with respect to the penalty parameter $k$ (Theorem 4.1). We believe that such a fixed-point perspective is novel in the literature, as it not only offers a constructive framework for characterizing solutions in the general case but also yields insights into how the penalty parameter affects the convergence rate.
@@ 1.2 Closely Related Literature
Our general formulation is largely inspired by Garg et al. [29], which investigates the SCSBP with the KL divergence as the penalty function $G$. Within that framework, the authors established the asymptotic convergence of the optimal policies as the penalty parameter $k$ tends to infinity, under the assumption that the initial measure is a delta measure. However, the use of KL divergence presents practical difficulties: if the model distribution $\mu$ assigns zero probability to any region where the data distribution $\mu_{\rm tar}$ has positive mass, then $\mathrm{KL}(\mu_{\rm tar}\|\mu)=+\infty$, rendering the divergence ill-posed under support mismatch [15, 38]. Moreover, a delta initial measure is rarely employed in generative tasks, as it lacks the diversity and randomness required for effective training. Finally, no convergence rate is quantified therein.
@@ 2 Preliminary
$V(\mathbb{P}):=\inf_{\mathbb{Q}\in\mathscr{P}^{\prime}}D_{\rm KL}(\mathbb{Q}\|\mathbb{P}),$ (2.1)
where $\mathscr{P}^{\prime}\subset\mathscr{P}(\Omega)$ is a given admissible set, and $D_{\rm KL}(\,\cdot\,\|\,\cdot\,)$ is the so-called Kullback-Leibler Divergence or the Relative Entropy (cf. [39]), defined by
@@
We remark here that the KL-divergence $D_{\rm KL}(\,\cdot\,\|\,\cdot\,)$ can be easily extended to any $\sigma$-finite measures [^2] In this case $D_{\rm KL}(\mathbb{Q}\|\mathbb{P}):=\int\log\big(\frac{\mathbb{Q}(dx)}{\mathbb{P}(dx)}\big)\mathbb{Q}(dx)$, if $d\mathbb{Q}\ll d\mathbb{P}$.. In what follows, we shall focus on the case when $\mathbb{Q}=\mathbb{P}\,\circ\,X^{-1}$ for some $\mathbb{R}^{d}$-valued continuous process $X$ defined on the canonical space with some $\mathbb{P}\in\mathscr{P}(\Omega)$, such that $\mathbb{Q}_{0}\negthinspace:=\negthinspace\mathbb{P}\circ{X_{0}}^{-1}={\mu_{\rm ini}}$, $\mathbb{Q}_{T}\negthinspace:=\negthinspace\mathbb{P}\circ{X_{T}}^{-1}\negthinspace=\negthinspace{\mu_{\rm tar}}\negthinspace$, and denote $\mathscr{P}^{\prime}:=\mathscr{P}({\mu_{\rm ini}},{\mu_{\rm tar}})\subseteq\mathscr{P}(\Omega)$ be all such “path measures”.
We note that if $\mathbb{P}=\mathbb{P}^{0}$ is Wiener measure and $\mathbb{Q}$ is equivalent to $\mathbb{P}^{0}$, then by the Cameron-Martin-Girsanov theorem, there exists $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d\times d})$, such that
@@
is a $\mathbb{P}^{0}$-martingale, and $\widetilde{W}_{t}:=W_{t}-\int_{0}^{t}\alpha_{s}ds$, $t\in[0,T]$, is a $\mathbb{Q}$-Brownian motion. It can then be easily calculated that
@@
In light of (2.2), one can easily associate a SBP to a stochastic control problem (see, e.g., [14, §4.4], [17, §1] and [40, p36]). Consider, for example, a standard SDE on $(\Omega,{\cal F},\mathbb{P}^{0})$:
$dX_{t}=b(t,X_{t})dt+dW_{t},\quad X_{0}\sim{\mu_{\rm ini}}.$ (2.3)
Denote $\mathbb{P}=\mathbb{P}_{X}:=\mathbb{P}^{0}\circ X^{-1}\in\mathscr{P}(\Omega)$. Then we consider the following SBP:
@@
Similar to (2.2), we can recast (2.4) as the following stochastic control problem:
@@
where $\mathbb{Q}\in\mathscr{P}(\Omega)$ is such that $\frac{d\mathbb{Q}}{d\mathbb{P}}=\mathscr{E}(\alpha)$ for some $\alpha\in\mathcal{A}\subseteq\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T])$, under which the underlying controlled dynamics takes the form:
$dX^{\alpha}_{t}=[b(t,X^{\alpha}_{t})+\alpha_{t}]dt+d\widetilde{W}_{t},\quad\mathbb{Q}\circ(X_{0}^{\alpha})^{-1}={\mu_{\rm ini}},\quad\mathbb{Q}\circ(X_{T}^{\alpha})^{-1}={\mu_{\rm tar}},$ (2.6)
@@
It is rather tempting to apply the idea above to the so-called McKean-Vlasov SBP (MVSBP). Consider, for example, the following McKean-Vlasov SDE on $(\Omega,{\cal F},\mathbb{P}^{0})$:
$dX_{t}=b(t,X_{t},\mathbb{P}_{t})dt+dW_{t},\quad X_{0}\sim{\mu_{\rm ini}}.$
where, again, we denote $\mathbb{P}=\mathbb{P}^{0}\circ X^{-1}\in\mathscr{P}_{2}(\Omega)$, and $\mathbb{P}_{t}:=\mathbb{P}^{0}\circ X^{-1}_{t}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, $t\in[0,T]$. Similar to (2.4) we can define an SBP, and let us refer to it as an MVSBP. Again, by (2.2), we can recast such MVSBP as the following (weak form) stochastic control problem:
$V({\mu_{\rm ini}},{\mu_{\rm tar}}):=\inf_{\alpha}\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{T}\frac{1}{2}|\alpha_{t}|^{2}\operatorname{{\rm d}}t\Big],$ (2.7)
where $\mathbb{Q}\in\mathscr{P}(\Omega)$ is such that the underlying controlled process $X=X^{\alpha}$ satisfies, under $\mathbb{Q}$:
$dX_{t}=[\alpha_{t}+b(t,X_{t},\mathbb{P}_{t})]dt+d\widetilde{W}_{t},\quad X_{0}\sim{\mu_{\rm ini}},\quad X_{T}\sim{\mu_{\rm tar}},$ (2.8)
where $\widetilde{W}$ is a $\mathbb{Q}$-Brownian motion, and we can assume that $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T])$. However, by looking at (2.8) more closely we see that the problem (2.7) and (2.8) do not form a McKean-Vlasov control problem, since $\mathbb{P}_{t}\neq\mathbb{Q}\circ(X_{t})^{-1}$(!). Therefore, an MVSBP should be formulated more carefully so as to connect to an McKean-Vlasov stochastic control problem. ∎
@@
(i) $\frac{d\mathbb{Q}}{d\mathbb{P}}\Big|_{{\cal F}_{T}}=\mathscr{E}(\alpha)=\exp\big\{\int_{0}^{T}\alpha_{s}dW_{s}-\frac{1}{2}\int_{0}^{T}|\alpha_{s}|^{2}ds\big\}$, $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d\times d})$;
@@
$dX^{\alpha}_{t}=[b(t,X^{\alpha}_{t})+\sigma(t)\alpha_{t}]dt+\sigma(t)d\widetilde{W}_{t},\quad\mathbb{Q}\circ(X_{0}^{\alpha})^{-1}={\mu_{\rm ini}},$ (2.9)
@@
The coefficients $b:[0,T]\times\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ and $\sigma:[0,T]\mapsto\mathbb{R}^{d\times d}$ are given deterministic continuous functions, such that there exists $L>0$, it holds that
@@
A smooth function $G(\cdot)=G(\cdot;{\mu_{\rm tar}}):\mathscr{P}_{2}(\mathbb{R}^{d})\to[0,\infty)$ is called a smooth penalty function associated to ${\mu_{\rm tar}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ if: $G(\mu;{\mu_{\rm tar}})=0$ if and only if $\mu={\mu_{\rm tar}}$. ∎
@@
For $k\in\mathbb{N}$, find $\alpha^{k}\in{\cal A}$ such that $V^{k}:=J^{k}(\widehat{\alpha}^{k})=\inf_{\alpha\in{\cal A}}J^{k}(\alpha)$, where
@@
To cast this problem into our framework, we can take, for example, ${\mu_{\rm ini}}=\mathcal{N}(0,I)$ and ${\mu_{\rm tar}}=p_{\rm data}$ in the theoretical framework (or ${\mu_{\rm tar}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ in the practical implementation). Then the optimal control $\alpha^{*}$ of SCSBP leads to a controlled process $(X_{t}^{\alpha^{*}})_{0\leq t\leq T}$ that simulates the data output $X_{T}^{\alpha^{*}}$. Our key results (see Theorem 4.1 and Theorem 6.4 below) show that the terminal measure $\mathbb{Q}_{X_{T}^{\alpha^{*}}}$ is close to $\mu_{\rm tar}$}, when $k$ is sufficiently large. ∎
@@
In terms of our framework, we can consider (2.3) as a pre-trained model with the drift $b(t,x):=s_{\widehat{\theta}}(t,x)$ being a well-trained score function, and $\widehat{\theta}$ is the trained parameter. Note that, as the result of pre-training, the output measure $\mathbb{Q}_{X_{T}}$ is sufficiently close to some data distribution $\mu_{\rm tar}$}. We then introduce a fine-tuning procedure through a reference measure $p_{\rm ref}$ with density $\frac{\exp(r(x))}{\int_{\mathbb{R}^{d}}\exp(r(x))dx}$, where $r:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a given reward function satisfying $\int_{\mathbb{R}^{d}}\exp(r(x))dx<\infty$. Now replacing ${\mu_{\rm tar}}=p_{\rm ref}$, the optimal control $\alpha^{*}$ of the corresponding SCSBP can then serve as the fine-tuning score function; and consequently, the new drift term $b(t,X^{\alpha*}_{t})+\sigma(t)\alpha^{*}$ acts as a combined score function.
Clearly, in this application the penalty parameter $k$ should not be chosen too large; otherwise, the effect of the preference function may dominate the fidelity to the original data distribution. With an appropriately selected $k$, the resulting measure $\mathbb{Q}_{X^{\alpha^{*}}_{T}}$ not only reflects $p_{\rm data}$ but also integrates the reward function $r$. In contrast, we remark that the classic SBP (2.1) is not capable of handling this application as it has a pre-fixed target distribution. ∎
@@
Transfer learning is a machine learning approach where knowledge gained from a “source task" is reused to improve learning in a related but different “target task" [9, 61, 48]. In what follows we shall consider transfer learning in the context of data generation.

### v2

@@ 1 Introduction
The recent generative modeling literature has seen a surge in the use of Schrödinger bridges. In these applications, one typically starts or chooses some distribution $\mu_{\rm ini}$} that is easy to sample from, and tries to “learn” an unknown distribution $\mu_{\rm tar}$} of a given data set. By numerically approximating the solution to the Schrödinger bridge problem, one can generate unlimited samples (i.e., synthetic data points that resemble the original data set). One such algorithm is presented by De Bortoli et al. [20] and Vargas et al. [64], who approximate the iterative proportional fitting procedure (Deming-Stephan [21]), using score matching with neural networks and maximum likelihood, respectively. Concurrently, Wang et al. [66] proposed a two-stage method with an auxiliary bridge handling possible non-smooth $\mu_{\rm tar}$}. Some more recent developments include [13, 56, 49, 52, 67, 32, 63, 54, 62, 2] as well as developing optimal transport techniques for generative AI tasks [44, 5, 42, 68, 1].
@@
In this paper, in light of Garg et al. [29], we study a soft-constrained Schrödinger bridge problem (SCSBP). Mathematically, we consider a penalty function $G:\mathscr{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}_{+}$, satisfying $G(\mu;{\mu_{\rm tar}})=0$ if and only if $\mu={\mu_{\rm tar}}$, where $\mathscr{P}_{2}(\mathbb{R}^{d})$ denotes all the probability measures on $\mathbb{R}^{d}$ with finite second moment equipped with 2-Wasserstein metric, and ${\mu_{\rm tar}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ is some given “target” distribution. For each $k\in\mathbb{N}$ and a given initial distribution ${\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, we consider the following stochastic control problem with dynamics on a given probability space $(\Omega,{\cal F},\mathbb{P})$:
@@
where $\mathbb{P}_{X^{\alpha}_{T}}$ is the law of $X^{\alpha}_{T}$ under probability $\mathbb{P}$ and the control $\alpha$ is chosen from a square-integrable progressively measurable admissible control set $\cal A$}. The goal is to find $V^{k}:=\inf_{\alpha\in{\cal A}}J^{k}(\alpha)$ and optimal control $\alpha^{k}$, for each $k\in\mathbb{N}$, and study the limiting behavior of $\{\alpha^{k}\}$ and $\{V^{k}\}$. Clearly, the dependence of the terminal cost on $\mathbb{P}_{X^{\alpha}_{T}}$ renders this relaxed formulation a non-standard stochastic control problem, leading to a McKean-Vlasov type of stochastic control. In contrast to Garg et al. [29], which focuses on the case where the penalty $G$ is given by the KL divergence and $\mu_{\rm ini}$ is a delta measure, we investigate the problem under more general cost functions and initial distributions.
Compared to existing methods using SBP under hard constraints, there are several advantages to considering SCSBP. First, when the KL divergence between $\mu_{\rm tar}$} and $\mathbb{P}_{X^{\alpha}_{T}}$ is infinite, the Schrödinger bridge does not admit a solution, whereas SCSBP always does (cf. e.g., [29]). More importantly, the penalty parameter $k$ acts as a regularization factor, preventing the algorithm from overfitting to $\mu_{\rm tar}$, which is crucial for certain generative modeling tasks with limited data [30, 45]. In addition, SCSBP provides a more general framework in generative AI, with applications beyond pre-training in data generation, and can be applied to fine-tuning and transfer learning (see Examples 2.6 and 2.7).
@@ 1.1 Outline of the Main Results and Contributions
The similar results in the case of a general initial distribution is much more involved. Among other things, we establish and/or extend some recently observed stability results of the solutions to the SBP, as the foundation for a fixed-point argument. A key element in our argument is the continuous dependence (or stability) of a mapping that is well known in the (static) optimal transport literature. Specifically, fix ${\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ and consider any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, it is known (cf. e.g., [7]) that there exists a unique pair of $\sigma$-finite measures $(\nu_{0},\nu_{T}):={\cal T}(\mu)$ such that the measure $\pi$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ defined by
has the marginals $\mu_{\rm ini}$} and $\mu$, where $p(\cdot,\cdot;\cdot,\cdot)$ is the transition density of a given diffusion process. Then it turns out that the solution to the SCSBP is the fixed-point of the mapping $\Gamma:\mathscr{P}_{2}(\mathbb{R}^{d})\to\mathscr{P}_{2}(\mathbb{R}^{d})$, where
$\Gamma(\mu):={\operatorname*{arg\,min}}_{\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{d})}\left\{kG(\bar{\mu})+\mathbb{E}_{X\sim\bar{\mu}}[\log\rho^{\mu}(X)]\right\},$ (1.2)
where $\rho^{\mu}$ is the density of the $\sigma$-finite measure $\nu$ in (1.1) (which may not be a probability measure). The successful application of Schauder’s fixed-point theorem on the Wasserstein space relies on several key elements, in particular a crucial continuous dependence result of the mapping $\mu\mapsto\rho^{\mu}$, for which we introduce an auxiliary entropic optimal transport problem, and identify its solution to the measure $\pi$ in (1.1). By utilizing some important stability results of the corresponding Schrödinger potentials (Proposition 5.9), together with some arguments in the spirit of the converse of Scheffé’s theorem (i.e., the weak convergence in Prohorov metric vs. the convergence in densities) as well as the so-called $\Gamma$-convergence of the minimizers of the optimization problem (1.2), we are able to verify the required properties so the mapping $\Gamma$ has a fixed-point (Theorem 6.1). As a direct consequence, we establish that the optimal control of the SCSBP converges linearly with respect to the penalty parameter $k$ (Theorem 6.4). We believe that such a fixed-point perspective is novel in the literature, as it not only offers a constructive framework for characterizing solutions in the general case but also yields insights into how the penalty parameter affects the convergence rate.
@@ 1.2 Closely Related Literature
Our general formulation is largely inspired by Garg et al. [29], which investigates the SCSBP with the KL divergence as the penalty function $G$. Within that framework, the authors established the asymptotic convergence of the optimal policies as the penalty parameter $k$ tends to infinity, under the assumption that the initial measure is a delta measure. However, the use of KL divergence presents practical difficulties: if the model distribution $\mu$ assigns zero probability to any region where the data distribution $\mu_{\rm tar}$ has positive mass, then $D_{\mathrm{KL}}(\mu_{\rm tar}\|\mu)=+\infty$, rendering the divergence ill-posed under support mismatch [15, 38]. Moreover, a delta initial measure is rarely employed in generative tasks, as it lacks the diversity and randomness required for effective training. Finally, no convergence rate is quantified therein.
@@ 2 Preliminary
$V(\mathbb{P}):=\inf_{\mathbb{Q}\in\mathscr{P}({\mu_{\rm ini}},{\mu_{\rm tar}})}D_{\rm KL}(\mathbb{Q}\|\mathbb{P}),$ (2.1)
where $D_{\rm KL}(\,\cdot\,\|\,\cdot\,)$ is the Kullback-Leibler Divergence or the Relative Entropy (cf. [39])  [^2] We remark here that the KL divergence $D_{\rm KL}(\,\cdot\,\|\,\cdot\,)$ can be easily extended to any $\sigma$-finite measures. In this case $D_{\rm KL}(\mathbb{Q}\|\mathbb{P}):=\int\log\big(\frac{\mathbb{Q}(dx)}{\mathbb{P}(dx)}\big)\mathbb{Q}(dx)$, if $d\mathbb{Q}\ll d\mathbb{P}$., defined by
@@
We note that the admissible set $\mathscr{P}({\mu_{\rm ini}},{\mu_{\rm tar}})\subseteq\mathscr{P}(\Omega)$ in (2.1) is defined as the collection of all probability measures $\mathbb{Q}\in\mathscr{P}(\Omega)$, such that $\mathbb{Q}$ is the law of some continuous $\mathbb{R}^{d}$-valued continuous process $X$ defined on the canonical space with probability measure ${\bar{\mathbb{P}}}\in\mathscr{P}(\Omega)$ (i.e., $\mathbb{Q}=\bar{\mathbb{P}}\,\circ\,X^{-1}$) satisfying ${\mathbb{Q}}\circ{X_{0}}^{-1}={\mu_{\rm ini}}$ and ${\mathbb{Q}}\circ{X_{T}}^{-1}\negthinspace=\negthinspace{\mu_{\rm tar}}\negthinspace$. In other words, $\mathscr{P}({\mu_{\rm ini}},{\mu_{\rm tar}})$ contains all “path measures” with initial and terminal distributions being $\mu_{\rm ini}$} and $\mu_{\rm tar}$}, respectively.
We note that if $\mathbb{P}=\mathbb{P}^{0}$ is Wiener measure and $\mathbb{Q}\in\mathscr{P}(\Omega)$ is absolutely continuous with respect to $\mathbb{P}^{0}$, then by the Cameron-Martin-Girsanov theorem, there exists $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d})$, such that
@@
is a $\mathbb{P}^{0}$-martingale, and $\widetilde{W}_{t}:=W_{t}-\int_{0}^{t}\alpha_{s}ds$, $t\in[0,T]$, is a $\mathbb{Q}$-Brownian motion. In this case, one can then easily check that
@@
In light of (2.2), one can easily associate a SBP to a stochastic control problem (see, e.g., [14, §4.4], [17, §1] and [40, p36]). Consider a standard SDE on a canonical space $(\Omega,{\cal F},\mathbb{P})$, with $\mathbb{P}\in\mathscr{P}(\Omega)$:
$dX_{t}=b(t,X_{t})dt+dW_{t},\quad\mathbb{P}\circ(X_{0})^{-1}={\mu_{\rm ini}}.$ (2.3)
In what follows we shall denote, for $t\in[0,T]$, $\mathbb{P}_{X_{t}}:=\mathbb{P}\circ(X_{t}^{-1})\in\mathscr{P}_{2}(\mathbb{R}^{d})$.
Then we consider the following SBP:
@@
Similar to (2.2), we now recast (2.4) as the following stochastic control problem:
@@
where $\mathbb{Q}\in\mathscr{P}(\Omega)$ is such that $\frac{d\mathbb{Q}}{d\mathbb{P}}=\mathscr{E}(\alpha)$ for some $\alpha\in\mathcal{A}\subseteq\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d})$, under which the underlying controlled dynamics takes the form:
$dX^{\alpha}_{t}=[b(t,X^{\alpha}_{t})+\alpha_{t}]dt+d\widetilde{W}_{t},\quad\mathbb{Q}_{X_{0}^{\alpha}}={\mu_{\rm ini}},\quad\mathbb{Q}_{X_{T}^{\alpha}}={\mu_{\rm tar}},$ (2.6)
@@
It is rather tempting to apply the idea above to the so-called McKean-Vlasov SBP (MVSBP). Consider, for example, the following McKean-Vlasov SDE on $(\Omega,{\cal F},\mathbb{P})$:
$dX_{t}=b(t,X_{t},\mathbb{P}_{X_{t}})dt+dW_{t},\quad\mathbb{P}_{X_{0}}={\mu_{\rm ini}}.$
Similar to (2.4) we can define an SBP, which we shall refer to as an MVSBP. Again, by (2.2), we recast such MVSBP as the following (weak form) stochastic control problem:
$V({\mu_{\rm ini}},{\mu_{\rm tar}}):=\inf_{\alpha}\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{T}\frac{1}{2}|\alpha_{t}|^{2}\operatorname{{\rm d}}t\Big].$ (2.7)
Then, $\mathbb{Q}\in\mathscr{P}(\Omega)$ must be such that the underlying controlled process $X^{\alpha}$ becomes, under $\mathbb{Q}$:
$dX^{\alpha}_{t}=[\alpha_{t}+b(t,X^{\alpha}_{t},{\mathbb{P}_{X^{\alpha}_{t}}})]dt+d\widetilde{W}_{t},\quad\mathbb{Q}_{X^{\alpha}_{0}}={\mu_{\rm ini}},\quad\mathbb{Q}_{X^{\alpha}_{T}}={\mu_{\rm tar}},$ (2.8)
where $\widetilde{W}$ is a $\mathbb{Q}$-Brownian motion, and $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d})$. However, by looking at (2.8) more closely we see that the problem (2.7) and (2.8) do not form a McKean-Vlasov control problem, since $\mathbb{P}_{X_{t}}\neq\mathbb{Q}_{X_{t}}$(!). Therefore, an MVSBP should be formulated more carefully so as to connect to an McKean-Vlasov stochastic control problem. ∎
@@
(i) $\frac{d\mathbb{Q}}{d\mathbb{P}}\Big|_{{\cal F}_{T}}=\mathscr{E}(\alpha)=\exp\big\{\int_{0}^{T}\alpha_{s}dW_{s}-\frac{1}{2}\int_{0}^{T}|\alpha_{s}|^{2}ds\big\}$, $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d})$;
@@
$dX^{\alpha}_{t}=[b(t,X^{\alpha}_{t})+\sigma(t)\alpha_{t}]dt+\sigma(t)d\widetilde{W}_{t},\quad\mathbb{Q}_{X_{0}^{\alpha}}={\mu_{\rm ini}},$ (2.9)
@@
The coefficients $b:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\to\mathbb{R}^{d\times d}$ are given deterministic continuous functions, such that there exists $L>0$, it holds that
@@
A function $G(\cdot)=G(\cdot;{\mu_{\rm tar}}):\mathscr{P}_{2}(\mathbb{R}^{d})\to[0,\infty)$ is called a continuous penalty function associated to ${\mu_{\rm tar}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ if: (1) $G(\cdot)$ is continuous on $\mathscr{P}_{2}(\mathbb{R}^{d})$; and (2) $G(\mu;{\mu_{\rm tar}})=0$ if and only if $\mu={\mu_{\rm tar}}$. ∎
@@
For $k\in\mathbb{N}$, find $\widehat{\alpha}^{k}\in{\cal A}$ such that $V^{k}:=J^{k}(\widehat{\alpha}^{k})=\inf_{\alpha\in{\cal A}}J^{k}(\alpha)$, where
@@
To cast this problem into our framework, we can take, for example, ${\mu_{\rm ini}}=\mathcal{N}(0,I)$ and ${\mu_{\rm tar}}=p_{\rm data}$ in the theoretical framework (or ${\mu_{\rm tar}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ in the practical implementation). Then the optimal control $\widehat{\alpha}$ of SCSBP leads to a controlled process $(X_{t}^{\widehat{\alpha}})_{0\leq t\leq T}$ that simulates the data output $X_{T}^{\widehat{\alpha}}$. Our key results (see Theorem 4.1 and Theorem 6.4 below) show that the terminal measure $\mathbb{Q}_{X_{T}^{\widehat{\alpha}}}$ is close to $\mu_{\rm tar}$}, when $k$ is sufficiently large. ∎
@@
In terms of our framework, we can consider (2.3) as a pre-trained model with the drift $b(t,x):=s_{\widehat{\theta}}(t,x)$ being a well-trained score function, and $\widehat{\theta}$ is the trained parameter. Note that, as the result of pre-training, the output measure $\mathbb{Q}_{X_{T}}$ is sufficiently close to some data distribution $\mu_{\rm tar}$}. We then introduce a fine-tuning procedure through a reference measure $p_{\rm ref}$ with density $\frac{\exp(r(x))}{\int_{\mathbb{R}^{d}}\exp(r(x))dx}$, where $r:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a given reward function satisfying $\int_{\mathbb{R}^{d}}\exp(r(x))dx<\infty$. Now replacing ${\mu_{\rm tar}}=p_{\rm ref}$, the optimal control $\widehat{\alpha}$ of the corresponding SCSBP can then serve as the fine-tuning score function; and consequently, the new drift term $b(t,X^{\widehat{\alpha}}_{t})+\sigma(t)\widehat{\alpha}$ acts as a combined score function.
Clearly, in this application the penalty parameter $k$ should not be chosen too large; otherwise, the effect of the preference function may dominate the fidelity to the original data distribution. With an appropriately selected $k$, the resulting measure $\mathbb{Q}_{X^{\widehat{\alpha}}_{T}}$ not only reflects $p_{\rm data}$ but also integrates the reward function $r$. In contrast, we remark that the classic SBP (2.1) is not capable of handling this application as it has a pre-fixed target distribution. ∎
@@
Transfer learning is a machine learning approach where knowledge gained from a “source task” is reused to improve learning in a related but different “target task” [9, 61, 48]. In what follows we shall consider transfer learning in the context of data generation.